# SWE-Compass: Towards Unified Evaluation of Agentic Coding Abilities for Large Language Models

### v1

@@ 5 Conclusion
We introduced SWE-Compass, a unified, execution-grounded benchmark that systematizes evaluation of agentic coding capabilities across 8 task types, 8 programming scenarios, and 10 languages with reproducible environments and task-type–aligned decision rules. Our large-scale study with two representative agent workflows (SWE-Agent and Claude Code) and 10 contemporary models reveals a clear difficulty hierarchy: tasks with deterministic oracles (Configuration & Deployment, Code Understanding) are substantially easier; feature-level tasks (Performance Optimization/Enhancement/Refactoring/Bug Fixing) occupy a mid tier; and Test Case Generation and Implementation remain the most challenging. The two agents are complementary: Claude Code excels where signals are stable and localized, while SWE-Agent is more resilient on investigative, multi-file problems that benefit from iterative localization.
Language-wise and scenario-wise analyses highlight that performance is governed less by raw coding difficulty than by tooling determinism, diagnosability, and repository observability. Our failure attribution shows that Requirement Misinterpretation and Incomplete Solution & Side Effects account for the majority of errors, while pure technical knowledge gaps are comparatively rare. Together, these findings suggest that future gains will come from improving requirement grounding, repository-level localization, execution reliability, and hypothesis pruning rather than from isolated advances in code generation alone. By providing a broad, production-aligned, and reproducible foundation, SWE-Compass aims to catalyze holistic progress in agentic coding systems.
6 Future Work
@@ Appendix A Appendix
A.1 Claude Code: Parallel Tool-Calls System Prompt
@@
options = ClaudeCodeOptions(
  max_turns=150,
  cwd=Path(working_dir),
  permission_mode="bypassPermissions",
  append_system_prompt=PARALLEL_TOOL_CALLS_SYSTEM_PROMPT.strip()
A.2 Trajectory Failure Analysis Prompt
@@ A.4 Executor Details and Method-Specific Settings
  • Claude Code. max turns =150=150; permission_mode == bypassPermissions; system prompt appended to encourage parallel tool calls (Appx. A.1); may internally invoke a SubAgent; networking disabled.
@@
MODEL Scores on Different Language
@@
A.5 Results by Data Source
Table 5: Eval Score (Pass@1, %) by data source on SWE-Compass (offline). “Open” aggregates public benchmarks (SWE-bench-Verified, Live, Fork, Multilingual); “Selected” are curated from GitHub PRs/Issues; “Generated” are controlled-synthesis instances exclusive to SWE-Compass. AVG is the macro-average across data sources.
MODEL Scores on Different Data Source
Open Selected Generated AVG
SWE-Agent
Claude-Sonnet-4 22.4 26.3 41.4 31.8
Qwen3-Coder-480B-Instruct 16.5 23.6 35.4 27.2
Kimi-K2-Instruct 10.2 18.5 32.4 22.7
Gemini-2.5-Pro 9.9 15.8 34.5 22.4
GPT-4.1 11.9 16.4 30.5 21.4
Qwen3-Coder-30B-Instruct 9.4 18.8 27.7 20.7
Qwen3-235B-A22B-Instruct 9.7 16.4 25.1 18.8
Gemini-2.5-Flash 8.2 12.9 28.7 18.5
Deepseek-V3 6.3 12.8 24.7 16.5
SWE-agent-LM-32B 6.8 12.5 20.4 14.7
Claude Code
Claude-Sonnet-4 19.3 27.9 43.8 32.9
Qwen3-Coder-480B-Instruct 12.2 15.3 32.9 21.9
Qwen3-Coder-30B-Instruct 7.4 14.0 31.6 21.6
Qwen3-235B-A22B-Instruct 8.0 10.8 21.6 14.7
Deepseek-V3 4.3 7.5 14.6 9.8
@@ A.5 Results by Data Source
A.6 Analysis of Open-Source Benchmark Distributions
@@
A.7 Model link list
We evaluate our approach using a diverse set of state-of-the-art language models, including both closed-source and open-source models. The closed-source models include Claude-Sonnet-4-20250514[Anthropic, 2025], Gemini-2.5-Flash, Gemini-2.5-Pro[Gemini Team and Google, 2023], and GPT-4.1-2025-04-14[OpenAI et al., 2024]. For open-source models, we utilize Qwen3-Coder series[Team, 2025, Hui et al., 2024b], Kimi-K2-Instruct-0905[Team et al., 2025a], Deepseek-V3-0324[Liu et al., 2024], and SWE-agent-LM-32B[Yang et al., 2025a]. The complete list of models with their official links is provided in Table 6.
Table 6: Model List

### v2

@@ 5 Conclusion
We introduced SWE-Compass, a unified, execution-grounded benchmark that enables systematic evaluation of large language models across diverse software engineering tasks, scenarios, and languages. By integrating 2,000 verified instances derived from real-world GitHub repositories with reproducible execution environments, SWE-Compass provides comprehensive coverage of the software development lifecycle. Our large-scale experiments with ten state-of-the-art LLMs under two agentic frameworks reveal consistent hierarchies of task difficulty, language-specific variability, and dominant failure modes rooted in requirement misinterpretation and incomplete solutions. These findings highlight that future progress in automated software engineering depends less on isolated code generation improvements and more on enhancing requirement grounding, environment reliability, and reasoning consistency. SWE-Compass offers a rigorous, scalable, and reproducible foundation for advancing the next generation of robust, general-purpose coding agents.
6 Future Works
@@ Appendix A Appendix
A.1 Judge Prompt for Code Understanding Task
Judge Prompt for Code Understanding Task
Evaluate if the answer satisfies the question requirements using a natural
language explanation.
QUESTION:
{question_text}
REQUIREMENTS (checklist items):
{chr(10).join(checklist_text)}
{patch_section}
ANSWER:
{truncated_answer}
EVALUATION RULES:
1. Answer MUST use clear English explanations, NOT just code diffs,
or other types of content
2. Only give 1.0 when ALL checklist items are thoroughly satisfied
with clear explanations.
3. Score = (satisfied items) / (total items)
4. Penalize: code diffs without explanation, vague statements,
wrong info. Give 0.0 when the answer is just code diffs or
completely wrong.
JSON response:
  "reasoning": "Brief explanation of which items satisfied/unsatisfied and why",
  "score": ,
  "satisfied_items": ["item_id1", ...]
A.2 Claude Code: Parallel Tool-Calls System Prompt
@@
A.3 Trajectory Failure Analysis Prompt
@@ A.4 Executor Details and Method-Specific Settings
  • Claude Code. max turns =150=150; permission_mode == bypassPermissions; system prompt appended to encourage parallel tool calls (Appx. A.2); may internally invoke a SubAgent; networking disabled.
@@
MODEL Scores on Different Languages
@@
A.5 Analysis of Open-Source Benchmark Distributions
@@
A.6 Model link list
We evaluate our approach using a diverse set of state-of-the-art language models, including both closed-source and open-source models. The closed-source models include Claude-Sonnet-4-20250514 [Anthropic, 2025], Gemini-2.5-Flash, Gemini-2.5-Pro [Gemini Team and Google, 2023], and GPT-4.1-2025-04-14 [OpenAI et al., 2024]. For open-source models, we utilize Qwen3-Coder series [Team, 2025, Hui et al., 2024b], Kimi-K2-Instruct-0905 [Team et al., 2025a], Deepseek-V3-0324 [Liu et al., 2024d], and SWE-agent-LM-32B [Yang et al., 2025a]. The complete list of models with their official links is provided in Table 5.
Table 5: Model List.